5a. A company with a perpetual inventory system purchases 2,000 items of inventory for $50 per item. The goods are shipped to the company as FOB destination. The freight cost is $1,000. The company then sells 1,100 items for $70 per item. These goods are shipped FOB destination to the customer at a cost of $450. Assuming that these are the only events that take place during this accounting period, what is the cost of inventory on hand at the end of the accounting period?

  1. $46,000

  2. $45,550

  3. $24,000

  4. $45,000

5b. What is Cost of Goods Sold? Gross margin? Operating income?
Answer #1

Cost of goods purchased = Number of units purchased x Cost per unit

= 2,000 x 50

= $100,000

Since the goods are shipped to the company as FOB destination, freight cost of $1000 will be borne by the seller.

Ending inventory units = Number of units purchased - Number of units sold

= 2,000 - 1,100

= 900

Cost of ending inventory = Ending inventory units x Cost per unit

= 900 x 50

= $45,000

Fourth option is correct.

Sales = Number of units sold x Selling price per unit

= 1,100 x 70

= $77,000

Cost of goods sold = Number of units sold x Cost per unit

= 1,100 x 50

= $55,000

Gross margin = Sales - Cost of goods sold

= 77,000 - 55,000

= $22,000

Operating income = Gross margin - Freight out

= 22,000 - 450

= $21,550

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.